Find p-adic norm of |3/4|_2


1
Expert's answer
2021-05-26T14:11:41-0400

The p-adic norm of "x" :

"|x|_p=p^{-a}"

where "p" is a prime number, "a" is a unique integer.

Then:

"\frac{3}{4}=3\cdot2^{-2}"


"|\frac{3}{4}|_2=2^2=4"